# End-to-End Large Portfolio Optimization for Variance Minimization with Neural Networks through Covariance Cleaning

### v1

@@ Abstract
We develop a rotation-invariant neural network that provides the global minimum-variance portfolio by jointly learning how to lag-transform historical returns and how to regularise both the eigenvalues and the marginal volatilities of large equity covariance matrices. This explicit mathematical mapping offers clear interpretability of each module’s role, so the model cannot be regarded as a pure black-box. The architecture mirrors the analytical form of the global minimum-variance solution yet remains agnostic to dimension, so a single model can be calibrated on panels of a few hundred stocks and applied, without retraining, to one thousand US equities—a cross-sectional jump that demonstrates robust out-of-sample generalisation. The loss function is the future realized minimum portfolio variance and is optimized end-to-end on real daily returns. In out-of-sample tests from January 2000 to December 2024 the estimator delivers systematically lower realised volatility, smaller maximum drawdowns, and higher Sharpe ratios than the best analytical competitors, including state-of-the-art non-linear shrinkage. Furthermore, although the model is trained end-to-end to produce an unconstrained (long–short) minimum-variance portfolio, we show that its learned covariance representation can be used in general optimizers under long-only constraints with virtually no loss in its performance advantage over competing estimators. These gains persist when the strategy is executed under a highly realistic implementation framework that models market orders at the auctions, empirical slippage, exchange fees, and financing charges for leverage, and they remain stable during episodes of acute market stress.
@@ 1 Introduction
The recent development of powerful machine learning approaches suggests that their use in portfolio optimization can provide results compatible with or better than state-of-the-art academic and professional solutions formalized in terms of mathematical finance models. Machine learning approaches to portfolio optimization can present different schemes. The most widespread approaches are: a) predict-then-optimize, b) decision-focused learning, and c) end-to-end approach [4]. Recent studies illustrate the extent of Neural Network (NN) applications to portfolio optimisation, yet they also expose recurring limitations that motivate the present work. Mesquita et al. [5] couple a Heterogeneous AutoRegressive filter with a multilayer perceptron to obtain positive-definite covariance forecasts, but the experiment involves only five equities and omits a Global Minimum Variance (GMV) benchmark, so scalability and incremental value remain unclear. Burda and Schroeder [6] embed a GO-GARCH model within a Long-Short Term Memory (LSTM) network to generate weights for the 100 largest U.S. stocks; the design yields modest gains over equal-weight and DCC strategies and is susceptible to survivor bias because delisted firms are excluded ex post. Huang et al. [7] adopt a two-stage network that first screens and then weights equities, yet the screening step discards securities that delist during the backtest, thus leaking future information. Reis et al. [8] combine a CNN and an LSTM to forecast covariances for a 14-asset mix of equities and fixed-income ETFs; fixed-income dominance in the GMV solution limits the interpretability of the results, and further evaluation on larger universes or on purely equity portfolios would be necessary to assess the model’s broader applicability. Collectively, these contributions demonstrate the promise of machine learning but highlight four issues: small or static universes, sub-optimal or heterogeneous benchmarks, potential data leakage, and uncertain behaviour when the cross-sectional dimension approaches the sample length. Addressing these gaps is the central objective of the architecture introduced below.
@@
One of the earliest regularization techniques is eigenvalue clipping. Laloux et al. [9] showed that the bulk of the eigenvalue spectrum of empirical financial covariance matrices closely follows the Marchenko–Pastur distribution for random noise. Eigenvalue clipping therefore replaces all sample eigenvalues that lie within this noise bulk by their average value. Later, Ledoit and Wolf [10] introduced the Linear Shrinkage (LS) estimator, which constructs a convex combination of the sample covariance matrix and a diagonal target matrix. This estimator remains well–defined when the number of assets $n$ exceeds the number of observations $\Delta t$. Its optimality in minimizing MSE holds strictly under the assumption that the true covariance matrix is drawn from an inverse–Wishart distribution [12]. In the last decade, Ledoit and collaborators introduced the NonLinear Shrinkage (NLS) estimator as an optimal method for covariance regularization [13]. Deriving the analytical form of the Ledoit–Péché (LP) estimator is challenging because it requires detailed knowledge of the sample eigenvector distribution [14]. The Quantized Eigenvalues Sampling Transform (QuEST) delivers the most accurate numerical implementation of the LP estimator by discretizing the limiting spectral distribution [15], but it incurs substantial computational cost. To alleviate this burden, the direct kernel estimator [16] and the inverse–Wishart regularization approach [11] offer closed-form approximations that closely mimic the LP solution. More recently, Ledoit and Wolf proposed the Quadratic-Inverse Shrinkage (QIS) estimator [17], which is considered the state-of-the-art technique in financial covariance filtering by combining computational efficiency with high estimation accuracy.
@@
Although the theoretical framework for covariance filtering is well established, its practical implementation remains challenging due to the pronounced non-stationarity of financial covariances. Market regimes evolve continuously under shifting economic conditions and investor behavior, so a covariance matrix estimated from past returns may poorly reflect current asset dependencies. Moreover, NLS theory assumes that sampling noise is the sole source of discrepancy between the sample and true covariance matrices and relies on the thermodynamic limit in which the number of assets $n$ and observations $\Delta t$ both tend to infinity with fixed aspect ratio $q=n/\Delta t$. In practice, extending $\Delta t$ to satisfy this limit only exacerbates non-stationarity, as longer historical windows encompass multiple structural shifts. Finally, the optimality of the LP estimator depends on the existence of finite twelfth-order moments of asset returns [14], yet empirical evidence indicates that equity returns exhibit heavy tails with tail exponents near four, rendering higher-order moments effectively unbounded. Methods based on $k$-fold Cross-Validation (CV) have been proposed to relax the stringent assumptions of nonlinear shrinkage theory [18, 12, 19]. These CV estimators achieve accuracy comparable to more complex schemes while permitting straightforward adaptations to specific data features. For example, Tan et al. [20] introduced an exponential decay factor into the return series within the CV framework. By applying exponential weights, one can maintain a long calibration window—approaching the thermodynamic limit—while emphasizing the most recent observations. This modification partially resolves the variance–bias trade-off caused by non-stationarity and structural shifts in financial covariances.
@@
The Engle Dynamical Conditional Covariance (DCC) model [28] represents a principled attempt to accommodate non-stationarity by modeling time-varying correlations through GARCH-type dynamics. It specifies separate equations for the conditional variances and then characterizes the evolving correlations of the standardized returns via a time-varying correlation matrix. . In principle, this approach should adapt to shifting market regimes and capture persistence in asset comovements. In practice, however, DCC is computationally intensive, often suffers from convergence issues, and yields OOS portfolios whose risk–return profiles rarely surpass those obtained with simpler static filters [29]. In contrast, the Average Oracle (AO) method [30, 29] abandons time-varying adjustments altogether. It fixes a set of eigenvalues calibrated over an extended historical window and applies them uniformly to all subsequent covariance estimates. This static correction ignores short-term noise but captures the enduring structure of the market. Empirical tests show that AO consistently outperforms both DCC-based and other NLS techniques, suggesting that, under strong non-stationarity, long-term averages provide more reliable covariance estimates than complex dynamic models [29, 31].
@@ 2.2 Eigenvalue Cleaning for MSE Minimization
Let $\widehat{\bm{C}}_{\text{MLE}}$ be the MLE of the correlation matrix obtained from $\Delta t_{\text{in}}$ observations drawn from a multivariate distribution with covariance matrix $\bm{\Sigma}=\bm{D}\bm{C}\bm{D}$, with $\bm{D}=\text{diag}(\sigma_{1},\ldots,\sigma_{n})$ being the diagonal matrix of the population assets’ volatilities and $\bm{C}$ the population correlation matrix, and let $\widehat{\bm{C}}_{\text{MLE}}=\widehat{\bm{V}}\widehat{\bm{\Lambda}}\widehat{\bm{V}}^{\top}$ be the spectral decomposition of the MLE. A RIE replaces the noisy eigenvalues by a deterministic function $\bm{f}$ while keeping $\widehat{\bm{V}}$ unchanged:
$\widehat{\mathbf{\Xi}}_{f}:=\widehat{\mathbf{V}}\,\bm{f}(\widehat{\bm{C}}_{\text{MLE}}\mid q)\,\widehat{\mathbf{V}}^{\top}.$ (5)
@@
$\bm{f}(\widehat{\bm{C}}_{\text{MLE}}\mid q)\longrightarrow\mathrm{diag}(\bm{f}(\widehat{\bm{\lambda}}\mid q)).$ (6)
Notably, $\bm{f}$ is a vector-valued function, i.e., $\bm{f}:\;\mathbb{R}^{n+1}\;\longrightarrow\;\mathbb{R}^{n}$, such that
$\bm{f}(\widehat{\lambda}_{1},\dots,\widehat{\lambda}_{n}\mid q)=\left\{f_{1}(\widehat{\lambda}_{1},\dots,\widehat{\lambda}_{n}\mid q),\;\dots,\;f_{n}(\widehat{\lambda}_{1},\dots,\widehat{\lambda}_{n}\mid q)\right\},\quad\text{with}\quad f_{i}:\mathbb{R}^{n+1}\to\mathbb{R};$ (7)
which means that it is generally a non-separable map since each component may depend on the entire argument. Furtermore, $\bm{f}$ is permuational-equivariant; therefore, called $\tau\in\mathcal{S}_{n}$ a permuation operator, then
@@
Because it requires knowledge of the (unknown) population correlation matrix $\bm{C}$, the exact solution of Eq. 9 is called an oracle $\bm{f}^{\mathrm{OE}}(\widehat{\bm{C}},\bm{C})$ and is not implementable in practice.
@@
$f^{\mathrm{OE}}_{i}(\widehat{\bm{C}},\bm{C})=\sum_{j=1}^{n}\lambda_{j}\,\psi_{ij},\qquad\psi_{ij}:=\sum_{k=1}^{n}(\widehat{v}_{ki}v_{kj})^{2}.$ (10)
where $\psi_{ij}$ is the overlap between the $i$-th eigenvector of the empirical correlation matrix and the $j$-th eigenvector of the population correlation matrix. Given that $\bm{\Psi}$ is an overlap between two reference frames, it is trivial to prove that it is a bi-stochastic matrix, i.e., the sum of each row and each column is equal to one. From this property comes that the oracle eigenvalues can be viewed as a weighted average of the population eigenvalues, where the weights are given by the overlap between the empirical and population eigenvectors. For bulk indices $i$, the weights $\psi_{ij}$ are widely dispersed over the $j$ elements in the bulk, so that $f_{i}^{\mathrm{OE}}$ becomes an almost uniform average of many population eigenvalues within the bulk. In a rough approximation, eigenvalue cleaning replaces the noisy fine structure of the bulk by a single effective level, while leaving the well-isolated high eigenvalues nearly untouched.
@@ 3 Neural Networks for Portfolio Optimization
The NN architecture that we propose is organised to replicate the analytical workflow of the GMV estimator. Embedding the algebraic symmetries of the GMV solution into the network topology both limits the number of free parameters and promotes robust OOS performance in large universes.
@@
Fig. 1 highlights the three learnable modules of our architecture. First, the lag-transformation module (Model 1, Sec. 3.1), applies a learnable temporal weighting kernel and a soft-clip to the raw return matrix $\bm{R}$, yielding the transformed returns $\widetilde{\bm{R}}$. Different lags can receive unequal weights; recent observations may be stressed, older observations may be down-weighted, or the reverse, depending on what minimises the loss. The resulting sequence bifurcates into two parallel branches.
The upper branch standardises $\widetilde{\bm{R}}$ to $\widetilde{\bm{Z}}$ and passes it to the eigenvalue-cleaning module (Model 2, Sec. 3.2). This block learns a non-linear shrinkage of the sample correlation matrix $\widetilde{\bm{C}}=\widetilde{\bm{Z}}\widetilde{\bm{Z}}^{\top}/\Delta t_{\mathrm{in}}$, producing a cleaned inverse eigenvalue matrix $\bm{\Lambda}_{\mathrm{NN}}^{-1}$.
The lower branch processes the vector of marginal volatilities $\left\{\widetilde{\sigma}_{1},\ldots,\widetilde{\sigma}_{n}\right\}$ of $\widetilde{\bm{R}}$ through the marginal-volatility network (Model 3, Sec. 3.3). This network transforms each standard deviation into its optimal inverse, generating the diagonal matrix $\bm{D}_{\mathrm{NN}}^{-1}=\operatorname{diag}(\bm{\sigma}_{\mathrm{NN}}^{-1})$.
@@
$\bm{\Sigma}_{\mathrm{NN}}^{-1}=\bm{D}_{\mathrm{NN}}^{-1}\widetilde{\bm{V}}\bm{\Lambda}_{\mathrm{NN}}^{-1}\widetilde{\bm{V}}^{\top}\bm{D}_{\mathrm{NN}}^{-1},$ (11)
where $\widetilde{\bm{V}}$ contains the eigenvectors of $\widetilde{\bm{C}}$. Substituting $\bm{\Sigma}_{\mathrm{NN}}^{-1}$ into $\widehat{\bm{\Sigma}}^{-1}$ of Eq. (1) yields the portfolio weights.
The architecture is trained end-to-end by minimising the realised OOS variance of the resulting portfolio. The loss function is

### v2

@@ Abstract
We develop a rotation-invariant neural network that provides the global minimum-variance portfolio by jointly learning how to lag-transform historical returns and marginal volatilities and how to regularise the eigenvalues of large equity covariance matrices. This explicit mathematical mapping offers clear interpretability of each module’s role, so the model cannot be regarded as a pure black box. The architecture mirrors the analytical form of the global minimum-variance solution yet remains agnostic to dimension, so a single model can be calibrated on panels of a few hundred stocks and applied, without retraining, to one thousand US equities—a cross-sectional jump that indicates robust generalization capability. The loss function is the future short-term realized minimum portfolio variance and is optimized end-to-end on real daily returns. In out-of-sample tests from January 2000 to December 2024, the estimator delivers systematically lower realized volatility, smaller maximum drawdowns, and higher Sharpe ratios than the best analytical competitors, including state-of-the-art non-linear shrinkage, and these advantages persist across both short and long evaluation horizons despite the model’s training focus on minimizing short-term variance. Furthermore, although the model is trained end-to-end to produce an unconstrained (long-short) minimum-variance portfolio, we show that its learned covariance representation can be used in general optimizers under long-only constraints with virtually no loss in its performance advantage over competing estimators. These gains persist when the strategy is executed under a highly realistic implementation framework that models market orders at the auctions, empirical slippage, exchange fees, and financing charges for leverage, and they remain stable during episodes of acute market stress.
@@ 1 Introduction
The literature on machine learning–based portfolio optimization is vast and rapidly expanding [4]. Existing studies can be grouped into three main categories: span predict-then-optimize pipelines, end-to-end differentiable frameworks, and reinforcement-learning agents. Overall, machine learning–based portfolio optimization is highly challenging with several limitations still to be overcome. In fact, some studies employ small or static asset universes that limit evidence of scalability and robustness. Benchmarks are in some cases infrequent or sub-optimal, making it difficult to compare performance improvements. Some designs inadvertently introduce data leakage through look-ahead bias or survivorship bias. These challenges constitute the principal obstacles to the practical deployment of machine learning in portfolio optimization.
@@
One of the earliest covariance regularization techniques is eigenvalue clipping. Laloux et al. [5] showed that the bulk of the eigenvalue spectrum of empirical financial covariance matrices closely follows the Marchenko–Pastur distribution for random noise. Eigenvalue clipping therefore replaces all sample eigenvalues that lie within this noise bulk by their average value. Later, Ledoit and Wolf [6] introduced the Linear Shrinkage (LS) estimator, which constructs a convex combination of the sample covariance matrix and a diagonal target matrix. This estimator remains well–defined when the number of assets $n$ exceeds the number of observations $\Delta t$. Its optimality in minimizing MSE holds strictly under the assumption that the true covariance matrix is drawn from an inverse–Wishart distribution [8]. In the last decade, Ledoit and collaborators introduced the NonLinear Shrinkage (NLS) estimator as an optimal method for covariance regularization [9]. Deriving the analytical form of the Ledoit–Péché (LP) estimator is challenging because it requires detailed knowledge of the sample eigenvector distribution [10]. The Quantized Eigenvalues Sampling Transform (QuEST) delivers the most accurate numerical implementation of the LP estimator by discretizing the limiting spectral distribution [11], but it incurs substantial computational cost. To alleviate this burden, the direct kernel estimator [12] and the inverse–Wishart regularization approach [7] offer closed-form approximations that closely mimic the LP solution. More recently, Ledoit and Wolf proposed the Quadratic-Inverse Shrinkage (QIS) estimator [13], which is considered the state-of-the-art technique in financial covariance filtering by combining computational efficiency with high estimation accuracy.
@@
Although the theoretical framework for covariance filtering is well established, its practical implementation remains challenging due to the pronounced non-stationarity of financial covariances. Market regimes evolve continuously under shifting economic conditions and investor behavior, so a covariance matrix estimated from past returns may poorly reflect current asset dependencies. Moreover, NLS theory assumes that sampling noise is the sole source of discrepancy between the sample and true covariance matrices and relies on the thermodynamic limit in which the number of assets $n$ and observations $\Delta t$ both tend to infinity with fixed aspect ratio $q=n/\Delta t$. In practice, extending $\Delta t$ to satisfy this limit only exacerbates non-stationarity, as longer historical windows encompass multiple structural shifts. Finally, the optimality of the LP estimator depends on the existence of finite twelfth-order moments of asset returns [10], yet empirical evidence indicates that equity returns exhibit heavy tails with tail exponents near four, rendering higher-order moments effectively unbounded. Methods based on $k$-fold Cross-Validation (CV) have been proposed to relax the stringent assumptions of nonlinear shrinkage theory [14, 15, 16, 8, 17]. These CV estimators achieve accuracy comparable to more complex schemes while permitting straightforward adaptations to specific data features. For example, Tan et al. [18] introduced an exponential decay factor into the return series within the CV framework. By applying exponential weights, one can maintain a long calibration window—approaching the thermodynamic limit—while emphasizing the most recent observations. This modification partially resolves the variance–bias trade-off caused by non-stationarity and structural shifts in financial covariances.
@@
The Engle Dynamical Conditional Covariance (DCC) model [26] represents a principled attempt to accommodate non-stationarity by modeling time-varying correlations through GARCH-type dynamics. It specifies separate equations for the conditional variances and then characterizes the evolving correlations of the standardized returns via a time-varying correlation matrix. In principle, this approach should adapt to shifting market regimes and capture persistence in asset comovements. In practice, however, DCC is computationally intensive, often suffers from convergence issues, and yields OOS portfolios whose risk–return profiles rarely surpass those obtained with simpler static filters [27]. In contrast, the Average Oracle (AO) method [28, 27] abandons time-varying adjustments altogether. It fixes a set of eigenvalues calibrated over an extended historical window and applies them uniformly to all subsequent covariance estimates. This static correction ignores short-term noise but captures the enduring structure of the market. Empirical tests show that AO consistently outperforms both DCC-based and other NLS techniques, suggesting that, under strong non-stationarity, long-term averages provide more reliable covariance estimates than complex dynamic models [27, 29].
@@
1.1 Literature review
The literature on NN portfolio optimization spans three methodological strands: predict-then-optimise pipelines, end-to-end differentiable architectures, and Reinforcement-Learning (RL) agents. We review some representative studies from every strand, briefly highlighting some design choices.
In the predict-then-optimise strand, Mesquita et al.,[33] combine a Heterogeneous AutoRegressive filter with a multilayer perceptron to deliver positive-definite covariance forecasts. They apply their method to a portfolio of five equities without providing evidence of scalability. Reis et al.,[34] train a 3-D CNN–BiLSTM–attention model on a 14-asset mixed universe before retraining it on 260 exchange-traded funds, showing scalability of the method. However, the relative contribution of fixed-income ETFs to the unconstrained GMV allocations is not explicitly discussed. Huang et al.,[35] first screen equities with an AGC-CNN that fuses OHLC images, graph convolutions, and self-attention, then feeds the survivors to a GMV allocator, obtaining improved Sharpe and Sortino ratios on CRSP stocks (2012–2023). The study is performed on the universe of 128 survivor firms of the considered period.
End-to-end differentiable frameworks embed the optimiser within the network so that forecasting and allocation are learned jointly. Butler and Kwon,[36] place the Markowitz module inside the computational graph via implicit differentiation of the KKT system, achieving lower out-of-sample costs than those obtained with ordinary least squares. The model is trained on a 24-contract futures universe by tuning a single exponential moving-average covariance window and with a fixed risk-aversion coefficient. Zhang et al.,[37] train an LSTM to map price histories of four index ETFs directly to long-only weights, maximising the Sharpe ratio. The method withstands the early-2020 Covid-19 shock. The set of assets comprises a relatively small ETF set. Costa and Iyengar,[38] embed a distributionally robust mean–variance layer that back-propagates through investor-risk and ambiguity-set parameters, outperforming equal-weight and classical pipelines on 2000–2021 US equities. The sample spans 20 surviving S&P 500 stocks. Uysal et al.,[39] integrate a differentiable risk-budgeting layer and raise the Sharpe ratio from 0.79 for nominal risk-parity to 1.24 for a portfolio of seven US ETFs (2011–2021).
RL-based studies treat portfolio choice as a sequential-decision task. Chaouki et al.,[40] use Deep Deterministic Policy Gradients on stylised markets with analytically known optima, showing agents that recover no-trade bands and threshold rules. Jang and Seong,[41] fuse technical indicators and long-horizon correlations through 3-D CNNs and tensor decomposition to guide an RL agent over 29 Dow components (2008–2019), achieving higher Sharpe ratios and lower drawdowns. The study is performed assuming zero slippage and fixed fees while excluding delisted firms. Yan et al.,[42] unite CNN and LSTM encoders with deterministic policy gradients to beat mean-reversion and moving-average baselines on cryptocurrencies and Chinese A-shares. Yu et al.,[43] embed a mean–variance layer inside an RL agent that reweights discrete-Fourier return components, outperforming equally weighted portfolio and static Markowitz on 47 CSI-300 survivors (2014–2024).
Across the differnt strands, three areas of improvement emerge: (i) asset universes should be increased to make applicability more realistic, dampening estimation risk, (ii) benchmark selection should be expanded and made less idiosyncratic to promote like-for-like comparisons, and (iii) careful planning of experimental protocols to avoide leakage of future information or understate implementation costs. We address these issues by training and testing on the full high- and mid-cap US equity universe, benchmarking against state-of-the-art minimum-variance strategies, and enforcing look-ahead-free splits with realistic frictions, so that observed performance gains reflect genuine model capability rather than experimental artefacts.
@@ 2.2 Eigenvalue Cleaning for MSE Minimization
Let us first consider the case when the covariance estimator is taken as $\widehat{\bm{C}}=\bm{C}_{\text{MLE}}$. Here $\bm{C}_{\text{MLE}}$ is the MLE of the correlation matrix obtained from $\Delta t_{\text{in}}$ observations drawn from a multivariate distribution with covariance matrix $\bm{\Sigma}=\bm{D}\bm{C}\bm{D}$, with $\bm{D}=\text{diag}(\sigma_{1},\ldots,\sigma_{n})$ being the diagonal matrix of the population assets’ volatilities and $\bm{C}$ the population correlation matrix, and let $\bm{C}_{\text{MLE}}=\widehat{\bm{V}}\widehat{\bm{\Lambda}}\widehat{\bm{V}}^{\top}$ be the spectral decomposition of the MLE. In what follows, we adopt the convention that a hat indicates quantities estimated via the MLE from sample data if not otherwise specified. A RIE replaces the noisy eigenvalues with a deterministic function $\bm{f}$ while keeping $\widehat{\bm{V}}$ unchanged:
$\mathbf{\Xi}_{\bm{f}}:=\widehat{\mathbf{V}}\,\bm{f}(\bm{C}_{\text{MLE}}\mid q)\,\widehat{\mathbf{V}}^{\top}.$ (5)
@@
$\bm{f}(\bm{C}_{\text{MLE}}\mid q)\mapsto\mathrm{diag}(\bm{f}(\widehat{\bm{\lambda}}\mid q)).$ (6)
Notably, $\bm{f}$ is a vector-valued function, i.e., $\bm{f}:\;\mathbb{R}^{n+1}\;\rightarrow\;\mathbb{R}^{n}$, such that
$\bm{f}(\widehat{\lambda}_{1},\dots,\widehat{\lambda}_{n}\mid q)=\left\{f_{1}(\widehat{\lambda}_{1},\dots,\widehat{\lambda}_{n}\mid q),\;\dots,\;f_{n}(\widehat{\lambda}_{1},\dots,\widehat{\lambda}_{n}\mid q)\right\},\quad\text{with}\quad f_{k}:\mathbb{R}^{n+1}\to\mathbb{R};$ (7)
which means that it is generally a non-separable map since each component may depend on the entire argument. Furthermore, $\bm{f}$ is permuational-equivariant; therefore, called $\tau\in\mathcal{S}_{n}$ a permutation operator, then
@@
Because it requires knowledge of the (unknown) population correlation matrix $\bm{C}$, the exact solution of Eq. (9) is called an oracle $\bm{f}^{\mathrm{OE}}(\widehat{\bm{C}},\bm{C})$ and is not implementable in practice.
@@
$f^{\mathrm{OE}}_{k}(\widehat{\bm{C}},\bm{C})=\sum_{l=1}^{n}\lambda_{l}\,\psi_{kl},\qquad\psi_{kl}:=\sum_{i=1}^{n}(\widehat{v}_{ik}v_{il})^{2}.$ (10)
where $\psi_{kl}$ is the overlap between the $k$-th eigenvector of the empirical correlation matrix and the $l$-th eigenvector of the population correlation matrix. Given that $\bm{\Psi}$ is an overlap between two reference frames, it is trivial to prove that it is a bi-stochastic matrix, i.e., the sum of each row and each column is equal to one. From this property comes that the oracle eigenvalues can be viewed as a weighted average of the population eigenvalues, where the weights are given by the overlap between the empirical and population eigenvectors. For bulk indices $k$, the weights $\psi_{kl}$ are widely dispersed over the $l$ elements in the bulk, so that $f_{k}^{\mathrm{OE}}$ becomes an almost uniform average of many population eigenvalues within the bulk. In a rough approximation, eigenvalue cleaning replaces the noisy fine structure of the bulk by a single effective level, while leaving the well-isolated high eigenvalues nearly untouched.
@@ 3 Neural Networks for Portfolio Optimization
The NN architecture that we propose is organized to replicate the analytical workflow of the GMV estimator. Embedding the algebraic symmetries of the GMV solution into the network topology both limits the number of free parameters and promotes robust OOS performance in large universes.
@@
Fig. 1 highlights the three learnable modules of our architecture. First, the lag-transformation module (Model 1, Sec. 3.1), applies a learnable temporal weighting kernel and a soft-clip to the raw return matrix $\bm{R}$, yielding the transformed returns $\widetilde{\bm{R}}$. Different lags can receive unequal weights; recent observations may be stressed, older observations may be down-weighted, or the reverse, depending on what minimizes the loss. The resulting sequence bifurcates into two parallel branches.
The lower branch standardizes $\widetilde{\bm{R}}$ to $\widetilde{\bm{Z}}$ and passes it to the eigenvalue-cleaning module (Model 2, Sec. 3.2). This block learns a non-linear shrinkage of the sample correlation matrix $\widetilde{\bm{C}}=\widetilde{\bm{Z}}\widetilde{\bm{Z}}^{\top}/\Delta t_{\mathrm{in}}$, producing a cleaned inverse eigenvalue matrix $\bm{\Lambda}_{\mathrm{NN}}^{-1}$.
The upper branch processes the vector of marginal volatilities $\left\{\widetilde{\sigma}_{1},\ldots,\widetilde{\sigma}_{n}\right\}$ of $\widetilde{\bm{R}}$ through the marginal-volatility network (Model 3, Sec. 3.3). This network transforms each standard deviation into its optimal inverse, generating the diagonal matrix $\bm{D}_{\mathrm{NN}}^{-1}=\operatorname{Diag}(\bm{\sigma}_{\mathrm{NN}}^{-1})$.
@@
$\bm{\Sigma}_{\mathrm{NN}}^{-1}=\bm{D}_{\mathrm{NN}}^{-1}\bm{V}_{\mathrm{NN}}\bm{\Lambda}_{\mathrm{NN}}^{-1}\bm{V}_{\mathrm{NN}}^{\top}\bm{D}_{\mathrm{NN}}^{-1}.$ (11)
The matrix $\bm{V}_{\mathrm{NN}}$ is obtained from the raw eigenvectors $\widetilde{\bm{V}}$ of the lag-transformed correlation $\widetilde{\bm{C}}$ and projected so that the estimated correlation has ones on its diagonal. In closed form,
$\bm{V}_{\mathrm{NN}}=\operatorname{Diag}\left(\operatorname{diag}\left(\widetilde{\bm{V}}\bm{\Lambda}_{\mathrm{NN}}\widetilde{\bm{V}}^{\top}\right)\right)^{-\frac{1}{2}}\widetilde{\bm{V}},$ (12)
where $\mathrm{diag}(\cdot)$ extracts the diagonal vector and $\mathrm{Diag}(\cdot)$ forms a diagonal matrix from that vector. Finally, substituting $\bm{\Sigma}_{\mathrm{NN}}^{-1}$ into $\widehat{\bm{\Sigma}}^{-1}$ of Eq. (1) yields the portfolio weights.
The architecture is trained end-to-end by minimizing the realized OOS variance of the resulting portfolio. The loss function is